OpenStudy (anonymous):

What is 15.27 as a mix fraction?

OpenStudy (anonymous):

15.27 = 1527/100 = (1500+27)/100 = 15 27/100

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

So do you always divide the decimal by 100 in that case

OpenStudy (anonymous):

there are two digits after he decimal so divide by 100 in case of three digits after decimal divide by 1000 and so on